what is x and y? answered with systems of equations solving with substitution
-8x-7y=-12, -7x-6y=-11

Answer by checkley77(12844)

-8x-7y=-12 multiply by 6
-7x-6y=-11 multioly by -7 & add.
-48x-42y=-72
49x+42y=77
----------------
x=5 ans.
-8*5-7y=-12
-40-7y=-12
-7y=-12+40
-7y=28
y=28/-7
y=-4 ans.
Proof:
-7*5-6*-4=-11
-35+24=-11
-11=-11